derive uncertainty principle using gaussian wave packet
1
Expert's answer
2020-01-22T04:23:37-0500

It is known that gaussian wave packet happen to be minimum uncertainly wave packets. In this case, if the uncertainty of coordinates Δx and the uncertainty of wave number Δk are taken as the standard deviations then this minimum value is "\frac{1}{2}."

In this way, we can write

"ΔxΔk≥\frac{1}{2} (1)"

Δx is the uncertainty of coordinates, Δk is the uncertainty of wave number


The wave number is equal to

"k=\frac{2πp}{h} (2)"

where h is the Planck constant, p is the momentum


Using (2) we have

"Δk=\frac{2πΔp}{h} (3)"

Where Δp is the uncertainty of momentum


We put (3) in (1)

"ΔxΔp≥\frac{ħ}{2}"

where "ħ=\frac{h}{2π}"